OpenStudy (ttop0816):

Evaluate cos7pi/6. -squareroot3/2 -1/2 1/2 squareroot3/2 <<< am i correct

OpenStudy (luigi0210):

Try again.

OpenStudy (luigi0210):

\[\LARGE cos\frac{7\pi}{6}=210^○\] Use the unit circle. http://i.stack.imgur.com/r8uHr.gif And remember. \[\LARGE (x, y)=(cos\theta, sin\theta)\]

OpenStudy (luigi0210):

You were off by a negative.

OpenStudy (ttop0816):

@Luigi0210 oh ): thank you!

OpenStudy (luigi0210):

Anytime.